# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ 4.1 Algorithm 1: Off-C2LUB
Compared to the online clustering of bandits problem, the core challenge in Off-ClusBand lies in data insufficiency. In the online setting, algorithms start with a complete graph over all users, where each node represents a user and each edge encodes a tentative assumption that the connected users are thought to be in the same cluster (Gentile et al., 2014; Li et al., 2019) and iteratively delete edges to separate users into clusters (i.e., remaining connected subgraphs). With abundant online data, this process accurately removes incorrect connecting edges. However, in the offline setting, limited data makes it difficult to precisely remove edges, often leading to the inclusion of heterogeneous users within the same cluster, which introduces bias and reduces decision quality. Therefore, building on the challenges described above, the core idea behind Off-C2LUB is to eliminate connections between users with significantly different preference vectors, i.e., prevent them from being assigned as neighbors. Specifically, the algorithm treats users with few samples independently, while leveraging the clustering property only for those with a sufficient dataset $\mathcal{D}_{u}$. The workflow of the proposed Off-C2LUB successfully avoids learning incorrect clusters by ensuring that predictions are made solely using data from users who share the same (or similar) preference vector as $u_{\textnormal{test}}$. The pseudo-code for Off-C2LUB, consisting a Cluster Phase and a Decision Phase, is presented in Algorithm 1.
@@
1:   Input: User $u_{\textnormal{test}}$, action set $A_{\textnormal{test}}$, dataset $\mathcal{D}$, parameters $\alpha>1$, $\lambda>0$, $\delta>0$ and $\hat{\gamma}\geq 0$
2:   Initialization: Construct a null graph $\mathcal{G}=(\mathcal{V},\emptyset)$, and compute $M_{u}$, $\bm{b}_{u}$, $\text{CI}_{u}$ as in Equation 1 and $\hat{\bm{\theta}}_{u}=M_{u}^{-1}\bm{b}_{u}$ for each user $u\in\mathcal{V}$
@@
5:      Connect $(u_{1},u_{2})$ if conditions in Equation 2 are satisfied
@@
7:  Let $\mathcal{G}_{\hat{\gamma}}=(\mathcal{V},\mathcal{E}_{\hat{\gamma}})$ denote the graph afterwards
@@
$\textstyle\tilde{\mathcal{D}}_{u}=\bigcup_{(u,v)\in\mathcal{E}_{\hat{\gamma}}}\left\{(\bm{a}_{v}^{i},r_{v}^{i})\in\mathcal{D}_{v}\right\}\cup\mathcal{D}_{u},\ \tilde{n}_{u}=1+\sum_{v\in\mathcal{V}}\mathbb{I}\left[(u,v)\in\mathcal{E}_{\hat{\gamma}}\right]$
$\textstyle\tilde{M}_{u}=\lambda\tilde{n}_{u}\cdot I+\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}\bm{a}\bm{a}^{\top},\ \tilde{N}_{u}=\left|\tilde{\mathcal{D}}_{u}\right|,\ \tilde{\bm{b}}_{u}=\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}r\bm{a},\ \tilde{\bm{\theta}}_{u}=\left(\tilde{M}_{u}\right)^{-1}\tilde{\bm{b}}_{u}$
@@
12:  Select $\bm{a}_{\textnormal{test}}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in A_{\textnormal{test}}}\left(\tilde{\bm{\theta}}_{u_{\textnormal{test}}}^{\top}\bm{a}-\beta\left\|\bm{a}\right\|_{\tilde{M}_{u_{\textnormal{test}}}^{-1}}\right)$ with $\beta=\sqrt{d\log\left(1+\frac{\tilde{N}_{u_{\textnormal{test}}}}{\lambda\tilde{n}_{u_{\textnormal{test}}}\cdot d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}$
Input and Initialization. Algorithm 1 takes as input the test user $u_{\textnormal{test}}\in\mathcal{U}$, the action set $A_{\textnormal{test}}$, the dataset $\mathcal{D}$ and several parameters (Line 1), which will be discussed later. The algorithm initializes the following (Line 2) for each user $u$:
$M_{u}=\lambda I+\sum_{i=1}^{N_{u}}\bm{a}_{u}^{i}\left(\bm{a}_{u}^{i}\right)^{\top},\ \bm{b}_{u}=\sum_{i=1}^{N_{u}}r_{u}^{i}\bm{a}_{u}^{i},\,\text{CI}_{u}=\frac{\sqrt{d\log\left(1+\frac{N_{u}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda_{a}N_{u}/2}}.$ (1)
Here, $M_{u}$ represents Gramian matrix regularized by the input parameter $\lambda$, $\bm{b}_{u}$ is the moment matrix of the response variable, $\hat{\bm{\theta}}_{u}$ is the ridge regression estimate of $\bm{\theta}_{u}$ for user $u$, and $\text{CI}_{u}$ is the confidence interval bound. These computed values are critical for both phases of the algorithm.
@@
$\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}<\hat{\gamma}-\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right)\text{ and }\min\left\{N_{u_{1}},N_{u_{2}}\right\}\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8Ud}{\lambda_{a}^{2}\delta}\right).$ (2)
The condition in Equation 2 ensures that the difference between the true preference vectors $\bm{\theta}_{u_{1}}$ and $\bm{\theta}_{u_{2}}$ is no greater than $\hat{\gamma}$ (see Theorem 4.2 for a detailed explanation). By only connecting users with similar estimated preference vectors, the algorithm prevents the introduction of bias that arises when users with highly divergent preference vectors are clustered together. Therefore, the choice of $\hat{\gamma}$ significantly affects the algorithm’s performance. As discussed in Section 4.2, a smaller $\hat{\gamma}$ helps avoid using samples from heterogeneous users of $u_{\text{test}}$, reducing bias, but may also limit access to samples from homogeneous users. In contrast, a larger $\hat{\gamma}$ includes more homogeneous samples to reduce variance, but risks introducing bias from heterogeneous users. We further discuss strategies for selecting $\hat{\gamma}$ in Section 4.3. Once all users in $\mathcal{V}$ have been checked, the graph is updated to $\mathcal{G}_{\hat{\gamma}}$, reflecting these new connections.
Next, for each user $u$, the algorithm aggregates data from the user and its neighbors in $\mathcal{G}_{\hat{\gamma}}$ to form a new dataset $\tilde{\mathcal{D}}_{u}$ (Line 9), which represents the samples from users who are likely to have the same preference vector as the test user. The algorithm then computes the necessary statistics for the decision phase. These statistics, similar to those defined in Equation 1, encompass a broader set of users with similar preferences as $u_{\text{test}}$, enhancing the algorithm’s decision-making capability.
Decision Phase. Finally, the algorithm selects the action $\bm{a}_{\textnormal{test}}$ for the test user $u_{\textnormal{test}}$ based on the computed statistics $\tilde{M}_{u_{\textnormal{test}}}$ and $\tilde{\bm{\theta}}_{u_{\textnormal{test}}}$ (Line 13), following a pessimistic estimate of the preference vector $\bm{\theta}_{u_{\textnormal{test}}}$ (Jin et al., 2021; Rashidinejad et al., 2021; Li et al., 2022). This approach ensures robust decision-making with offline data.
@@ 4.2 Upper Bound for Off-C2LUB
We analyze the upper bound for Algorithm 1’s suboptimality. Detailed proofs can be found in Appendices A.1. We denote the total number of samples for a set of users $\mathcal{S}$ as $N^{\mathcal{S}}=\sum_{v\in\mathcal{S}}N_{v}$. Since our analysis focuses on the suboptimality for the test user $u_{\text{test}}$, we refer to homogeneous neighbors as those neighbors of $u_{\text{test}}$ in the final graph $\mathcal{G}_{\hat{\gamma}}$ that belong to the same cluster, and to heterogeneous neighbors as those in different clusters. Moreover, in order to formalize the separation between clusters, we define the minimum heterogeneity gap between users from different clusters (Gentile et al., 2014, 2017; Wang et al., 2025):
Definition 4.1 (Minimum Heterogeneity Gap).
The preference vectors of users from different clusters are separated by at least a gap of $\gamma$. Specifically, two users $u$ and $v$ belong to the same cluster if and only if $\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}<\gamma$.
@@
Let $K_{u}=\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\lambda_{a}N_{u}}}$, which represents the extent of data insufficiency for user $u$. Then for given inputs $\alpha\geq 1$, $\lambda$ and $\delta$ such that $\lambda\leq d\log\left(1+\min_{u}\{N_{u}\}/(\lambda d)\right)+2\log\left(2U/\delta\right)$ and $\delta\leq\frac{2U}{1+\max_{u}\{N_{u}\}/(\lambda d)}$, the suboptimality gap for Off-C2LUB satisfies:
[…]
with probability at least $1-2\delta$, where $\tilde{O}$ omits logarithmic terms and
$\mathcal{R}_{\hat{\gamma}}(u)$ =$=\begin{cases}\{u\}\cup\left\{v\,|\,\bm{\theta}_{u}=\bm{\theta}_{v}\text{ and }K_{u}+K_{v}\leq\alpha_{r}\hat{\gamma}\text{ and }N_{v}\geq N_{\min}\right\},&\text{if }N_{u}\geq N_{\min},\\ \{u\},&\text{otherwise},\end{cases}$
$\mathcal{W}_{\hat{\gamma}}(u)$ =$=\begin{cases}\left\{v\,|\,\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma}\text{ and }K_{u}+K_{v}\leq\alpha_{w}\varepsilon\text{ and }N_{v}\geq N_{\min}\right\},&\text{if }N_{u}\geq N_{\min},\\ \emptyset,&\text{otherwise},\end{cases}$
@@
$\mathcal{V}_{\hat{\gamma}}(u)$ $=\mathcal{R}_{\hat{\gamma}}(u)\cup\mathcal{W}_{\hat{\gamma}}(u),\ \eta_{\mathcal{W}_{\hat{\gamma}}(u)}=\frac{\left|\mathcal{W}_{\hat{\gamma}}(u)\right|\lambda+N^{\mathcal{W}_{\hat{\gamma}}(u)}}{|\mathcal{V}_{\hat{\gamma}}(u)|\lambda+N^{\mathcal{V}_{\hat{\gamma}}(u)}}$
for $N_{\min}=\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right),\alpha_{r}=\frac{1}{2(\alpha+1)\sqrt{\max\{2,d\}}}$, $\alpha_{w}=\frac{1}{\alpha-1}$ and $\varepsilon=\hat{\gamma}-\gamma$.
Remark 4.3 (Notation Interpretations).
In theoretical analysis, the exact neighbors of $u_{\text{test}}$ are difficult to determine due to limited access to the full dataset $\mathcal{D}$. To address this, Algorithm 1 estimates neighbor sets using conditions from Line 5. The set $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$ serves as a subset estimation of homogeneous neighbors, conservatively underestimating their number to ensure that the first term in Equation 3 safely bounds bias from limited homogeneous samples. Conversely, $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ is a superset estimation of heterogeneous neighbors, deliberately overestimated to ensure the second term captures worst-case heterogeneity via a larger $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}$. These constructions allow us to derive a high-probability upper bound that remains valid despite cluster uncertainty. According to Equation 2, connections are formed only when both users meet a minimal data requirement $N_{\min}$, ensuring sufficient coverage of $\mathbb{R}^{d}$ per Assumption 3.1. Otherwise, the algorithm falls back on using only test user data. The conditions on $K_{u}+K_{v}$ for both $\mathcal{R}_{\hat{\gamma}}$ and $\mathcal{W}_{\hat{\gamma}}$ reflect our principle that only users with enough data and reliable estimates should influence decisions. A larger $\hat{\gamma}$ relaxes these constraints, potentially adding useful homogeneous samples but increasing the risk of heterogeneous bias. Finally, the constraint $\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma}$ ensures $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ includes only similar users, with larger $\hat{\gamma}$ risking more bias.
Remark 4.4 (Analysis of Theorem 4.2).
The terms in Theorem 4.2 reflect both homogeneous error and heterogeneous error. The first term represents the homogeneous error associated with homogeneous neighbors in $\mathcal{R}_{\hat{\gamma}}(u_{\textnormal{test}})$. Specifically, this term decreases as the total number of samples in $\mathcal{R}_{\hat{\gamma}}(u_{\textnormal{test}})$ increases, which can be achieved by selecting a larger $\hat{\gamma}$. The second term represents the heterogeneous error, which is directly proportional to $\hat{\gamma}$. As discussed in the previous remark, increasing $\hat{\gamma}$ enlarges the heterogeneous neighbor set $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$ and increases $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})}$. Therefore, controlling $\hat{\gamma}$ introduces a trade-off between expanding $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$ to include more useful samples and limiting $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ to reduce bias in the estimation, and is critical to prevent large heterogeneous errors while still leveraging enough data for prediction to decrease homogeneous errors.

### v2

@@ 4.1 Algorithm 1: Off-C2LUB
Compared to the online clustering of bandits problem, the core challenge in Off-ClusBand lies in data insufficiency. In the online setting, algorithms start with a complete graph over all users, where each node represents a user and each edge encodes a tentative assumption that the connected users are thought to be in the same cluster (Gentile et al., 2014; Li and Zhang, 2018) and iteratively delete edges to separate users into clusters (i.e., remaining connected subgraphs). With abundant online data, this process accurately removes incorrect connecting edges. However, in the offline setting, limited data makes it difficult to precisely remove edges, often leading to the inclusion of heterogeneous users within the same cluster, which introduces bias and reduces decision quality. Therefore, building on the challenges described above, the core idea behind Off-C2LUB is to prevent users with significantly different preference vectors from being clustered. Specifically, the algorithm treats users with few samples independently, while leveraging the clustering property only for those with a sufficient dataset $\mathcal{D}_{u}$. The workflow of the proposed Off-C2LUB successfully avoids clustering largely heterogeneous users by ensuring that predictions are made solely using data from users who share the same (or similar) preference vector as $u_{\textnormal{test}}$. Algorithm 1 shows the pseudo-code for Off-C2LUB which consists a Cluster Phase and a Decision Phase.
@@
1:Input: User $u_{\textnormal{test}}$, action set $\mathcal{A}_{\textnormal{test}}$, dataset $\mathcal{D}$, parameters $\alpha>1$, $\lambda>0$, $\delta>0$ and $\hat{\gamma}\geq 0$
2:Initialization: Construct a null graph $\mathcal{G}=(\mathcal{V},\emptyset)$ where $\mathcal{V}=\mathcal{U}$, set $N_{\min}=\frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8Ud}{\tilde{\lambda}_{a}^{2}\delta}\right)$, and compute $M_{u}$, $\bm{b}_{u}$, $\text{CI}_{u}$ as in Equation˜1 and $\hat{\bm{\theta}}_{u}=M_{u}^{-1}\bm{b}_{u}$ for each user $u\in\mathcal{U}$
@@
5:   Connect $(u_{1},u_{2})$ if conditions in Equation˜2 are satisfied
@@
[…]
@@
[…]
$\textstyle\tilde{M}_{u}=\lambda\tilde{n}_{u}\cdot I+\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}\bm{a}\bm{a}^{\top},\ \tilde{N}_{u}=\big|\tilde{\mathcal{D}}_{u}\big|,\ \tilde{\bm{b}}_{u}=\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}r\bm{a},\ \tilde{\bm{\theta}}_{u}=\big(\tilde{M}_{u}\big)^{-1}\tilde{\bm{b}}_{u}$
@@
12: Select $\bm{a}_{\textnormal{test}}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in\mathcal{A}_{\textnormal{test}}}\left(\tilde{\bm{\theta}}_{u_{\textnormal{test}}}^{\top}\bm{a}-\beta\left\|\bm{a}\right\|_{\tilde{M}_{u_{\textnormal{test}}}^{-1}}\right)$, $\beta=\sqrt{d\log\left(1+\frac{\tilde{N}_{u_{\textnormal{test}}}}{\lambda\tilde{n}_{u_{\textnormal{test}}}\cdot d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}$
Input and Initialization. Algorithm 1 takes as input the test user $u_{\textnormal{test}}\in\mathcal{U}$, the action set $\mathcal{A}_{\textnormal{test}}$, the dataset $\mathcal{D}$ and several parameters (Line 1), which will be discussed later. The algorithm initializes the following (Line 2) for each user $u$:
$M_{u}=\lambda I+\sum_{i=1}^{N_{u}}\bm{a}_{u}^{i}\left(\bm{a}_{u}^{i}\right)^{\top},\ \bm{b}_{u}=\sum_{i=1}^{N_{u}}r_{u}^{i}\bm{a}_{u}^{i},\,\text{CI}_{u}=\frac{\sqrt{d\log\left(1+\frac{N_{u}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\tilde{\lambda}_{a}N_{u}/2}}.$ (1)
Here, $M_{u}$ denotes the Gramian matrix regularized by $\lambda$, $\bm{b}_{u}$ is the moment matrix of the response variable, $\hat{\bm{\theta}}_{u}$ is the ridge regression estimate of $\bm{\theta}_{u}$ for user $u$, and $\text{CI}_{u}$ represents the confidence interval bound. The algorithm also defines $N_{\min}$ as the minimum number of samples required for each user during clustering. These quantities play a central role in both phases of the algorithm.
@@
$\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}<\hat{\gamma}-\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right)\text{ and }\min\left\{N_{u_{1}},N_{u_{2}}\right\}\geq N_{\min}.$ (2)
The condition in Equation˜2 ensures that the difference between the true preference vectors $\bm{\theta}_{u_{1}}$ and $\bm{\theta}_{u_{2}}$ is no greater than $\hat{\gamma}$ (see Lemma˜4.2 for a detailed explanation). By only connecting users with similar estimated preference vectors, the algorithm prevents the introduction of bias that arises when users with highly divergent preference vectors are connected. Therefore, the choice of $\hat{\gamma}$ significantly affects the algorithm’s performance. As discussed in Section 4.2, a smaller $\hat{\gamma}$ helps to reduce bias from heterogeneous users of $u_{\text{test}}$, but may also limit access to samples from homogeneous users. In contrast, a larger $\hat{\gamma}$ includes more samples to reduce noise, but risks introducing bias from heterogeneous users. We further discuss strategies for selecting $\hat{\gamma}$ in Section 4.3. Once all users in $\mathcal{V}$ have been checked, the graph is updated to $\mathcal{G}_{\hat{\gamma}}$, reflecting these new connections.
Next, for each user $u$, the algorithm aggregates data from the user and its neighbors in $\mathcal{G}_{\hat{\gamma}}$ to form a new dataset $\tilde{\mathcal{D}}_{u}$ (Line 9), which represents the samples from users who are likely to have the same preference vector as the test user. Unlike traditional online algorithms that aggregate data from the entire connected component containing $u$, our approach aggregates only from one-hop neighbors. We retain the term “clustering” in our algorithm’s name to stay consistent with the problem setting and prior works. This conservative strategy reduces uncontrolled bias from highly heterogeneous users (Wang et al., 2023a; Dai et al., 2024a) and is better suited to offline scenarios where data is limited, as discussed in detail in Appendix F. The algorithm then computes the necessary statistics for the decision phase. These statistics, similar to those defined in Equation˜1, encompass a broader set of users with similar preferences as $u_{\text{test}}$, enhancing the algorithm’s decision-making capability.
Decision Phase. Finally, the algorithm selects $\bm{a}_{\textnormal{test}}$ for the test user $u_{\textnormal{test}}$ based on the statistics $\tilde{M}_{u_{\textnormal{test}}}$ and $\tilde{\bm{\theta}}_{u_{\textnormal{test}}}$ (Line 9), following a pessimistic estimate of the preference vector $\bm{\theta}_{u_{\textnormal{test}}}$ (Jin et al., 2021; Rashidinejad et al., 2021; Li et al., 2022), ensuring robust decision-making with offline data.
@@ 4.2 Upper Bound for Off-C2LUB
We analyze the upper bound on the suboptimality of Algorithm 1, with detailed proofs provided in Appendix C.1 and C.2. Since the algorithm aggregates user data based on the graph $\mathcal{G}_{\hat{\gamma}}$ constructed in Line 7, we define $\mathcal{V}_{\hat{\gamma}}(u),\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$ as Table 2.
Table 2: Summary of neighbor set notations.
Notation Definition Interpretation
$\mathcal{V}_{\hat{\gamma}}(u)$ $\{u\}\cup\{v\mid(u,v)\in\mathcal{E}_{\hat{\gamma}}\}$ Set of user $u$ together with all its neighbors in graph $\mathcal{G}_{\hat{\gamma}}$.
$\mathcal{R}_{\hat{\gamma}}(u)$ $\{v\mid v\in\mathcal{V}_{\hat{\gamma}}(u),\bm{\theta}_{u}=\bm{\theta}_{v}\}$
Set of $u$ and its homogeneous neighbors, i.e., users in $\mathcal{V}_{\hat{\gamma}}(u)$ with same preference vectors.
Their data can be safely aggregated with $u$’s without introducing bias.
$\mathcal{W}_{\hat{\gamma}}(u)$ $\{v\mid v\in\mathcal{V}_{\hat{\gamma}}(u),\bm{\theta}_{u}\neq\bm{\theta}_{v}\}$
Set of $u$’s heterogeneous neighbors, i.e., users in $\mathcal{V}_{\hat{\gamma}}(u)$ with different preference vectors.
Aggregating their data with $u$’s may introduce bias and thus should be controlled.
Thus $\mathcal{V}_{\hat{\gamma}}(u)=\mathcal{R}_{\hat{\gamma}}(u)\cup\mathcal{W}_{\hat{\gamma}}(u)$ and $\mathcal{R}_{\hat{\gamma}}(u)\cap\mathcal{W}_{\hat{\gamma}}(u)=\emptyset$ hold. To formalize cluster separation, we introduce the heterogeneity gap (Gentile et al., 2014; Li et al., 2019, 2025; Wang et al., 2025) as:
Definition 4.1 (Heterogeneity Gap).
The preference vectors of users from different clusters are separated by at least $\gamma$; that is, for any $u$ and $v$ in different clusters, $\left\|\bm{\theta}_{u}-\bm{\theta}_{v}\right\|_{2}\geq\gamma>0$.
In Algorithm 1, the choice of $\hat{\gamma}$ influences both $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$, and thus directly impacts both performance and theoretical guarantees. The following lemma formalizes these effects, while Theorem 4.4 establishes the resulting suboptimality bound.
Lemma 4.2 (Estimations of sets $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$).
For inputs $\alpha\geq 1$, $\lambda>0$, and $\delta\in(0,1)$ satisfying $\lambda\leq d\log\big(1+\frac{\min_{u}\{N_{u}\}}{\lambda d}\big)+2\log\big(\frac{2U}{\delta}\big)\text{ and }\delta\leq\frac{2U}{1+\max_{u}\{N_{u}\}/(\lambda d)},$ there exist some $\alpha_{r}\in\Big[\frac{\sqrt{\tilde{\lambda}_{a}}}{4(\alpha+1)\sqrt{\log(2U/\delta)\max\{2,d\}}},\frac{\sqrt{\tilde{\lambda}_{a}}}{2\alpha\sqrt{\log(2U/\delta)}}\Big)$ and $\alpha_{w}\in\Big(0,\frac{\sqrt{\tilde{\lambda}_{a}}}{2(\alpha-1)\sqrt{\log(2U/\delta)}}\Big]$ such that sets $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$ defined in Table˜2 can be represented as:
$\text{If }N_{u}\geq N_{\min}:$ $\mathcal{R}_{\hat{\gamma}}(u)=\{u\}\cup\left\{v\,\middle|\,\bm{\theta}_{u}=\bm{\theta}_{v},\;\tfrac{1}{\sqrt{N_{u}}}+\tfrac{1}{\sqrt{N_{v}}}\leq\alpha_{r}\hat{\gamma},\;N_{v}\geq N_{\min}\right\},$
$\mathcal{W}_{\hat{\gamma}}(u)=\left\{v\,\middle|\,\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma},\;\tfrac{1}{\sqrt{N_{u}}}+\tfrac{1}{\sqrt{N_{v}}}\leq\alpha_{w}\varepsilon,\;N_{v}\geq N_{\min}\right\};$
$\text{Otherwise}:$ $\mathcal{R}_{\hat{\gamma}}(u)=\{u\},\mathcal{W}_{\hat{\gamma}}(u)=\emptyset,$
with probability at least $1-\delta$, where $N_{\min}$ is as defined in Line 2 of Algorithm˜1 and $\varepsilon=\hat{\gamma}-\gamma$.
Remark 4.3 (Interpretation of Lemma˜4.2).
As defined in Table˜2, set $\mathcal{R}_{\hat{\gamma}}(u)$ contains user $u$ and its homogeneous neighbors while $\mathcal{W}_{\hat{\gamma}}(u)$ represents $u$’s heterogeneous neighbors. In Lemma˜4.2, the distinction between the two cases (whether $N_{u}\geq N_{\min}$) follows directly from the cluster condition in Equation˜2. Specifically, when $u$ does not meet the minimum data requirement $N_{\min}$, no connections are formed, and the algorithm defaults to using only $u$’s own data. This threshold ensures sufficient coverage of $\mathbb{R}^{d}$ under Assumption 3.1, providing essential information along each dimension for preference estimation. When $N_{u}\geq N_{\min}$, the constraints on $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}$ for both $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$ reflect the principle that only users with sufficient data and reliable estimates should contribute to decision-making. Increasing $\hat{\gamma}$ (and thus $\varepsilon$) relaxes these constraints, potentially incorporating more homogeneous samples but at the risk of introducing larger bias. Moreover, the constraint $\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma}$ ensures that $\mathcal{W}_{\hat{\gamma}}(u)$ only includes users with relatively similar preferences where $\hat{\gamma}$ is a tunable input. A larger $\hat{\gamma}$ increases such bias and enlarges $\mathcal{W}_{\hat{\gamma}}(u)$.
@@
With the same conditions as Lemma˜4.2, the suboptimality for Off-C2LUB satisfies:
@@
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-C}^{2}\textnormal{LUB},\mathcal{A}_{\textnormal{test}})\leq\tilde{O}\left(\sqrt{\frac{d}{\tilde{\lambda}_{a}N_{\mathcal{V}_{\hat{\gamma}}(u_{\textnormal{test}})}}}+\frac{\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})}\hat{\gamma}}{\tilde{\lambda}_{a}}\right),$ (3)
with probability at least $1-2\delta$, where $\tilde{O}$ omits logarithmic terms and $\eta_{\mathcal{W}_{\hat{\gamma}}(u)}=\frac{\left|\mathcal{W}_{\hat{\gamma}}(u)\right|\lambda+N_{\mathcal{W}_{\hat{\gamma}}(u)}}{|\mathcal{V}_{\hat{\gamma}}(u)|\lambda+N_{\mathcal{V}_{\hat{\gamma}}(u)}}$ denotes the fraction of heterogeneous samples over all utilized samples for $u$.
Remark 4.5 (Analysis of Theorem˜4.4).
The terms in Theorem˜4.4 reflect both noise and bias. The first term, caused by the limited number of aggregated samples, represents the noise from all neighbors of $u_{\text{test}}$ in $\mathcal{V}_{\hat{\gamma}}(u_{\textnormal{test}})$. This noise decreases as the total number of samples in $\mathcal{V}_{\hat{\gamma}}(u_{\textnormal{test}})$ grows, which can be achieved by selecting a larger $\hat{\gamma}$. The second term reflects the bias from including samples of heterogeneous neighbors, which scales with both $\hat{\gamma}$ and the fraction of heterogeneous samples $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}$. As noted in Remark˜4.3, enlarging $\hat{\gamma}$ expands $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$ and thus potentially increases $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})}$. Therefore, controlling $\hat{\gamma}$ is key to balancing two competing effects: expanding $\mathcal{V}_{\hat{\gamma}}(u_{\text{test}})$ to reduce noise, and restricting $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$ to limit bias. The dependence on $\tilde{\lambda}_{a}$ highlights the link between suboptimality and action regularity: a larger $\tilde{\lambda}_{a}$ implies more information per sample and hence smaller suboptimality. Importantly, the form of Theorem˜4.4 reflects the fundamental limited-data challenge of the offline setting: unlike online scenarios, where additional samples can be collected to reduce noise and accurately estimate user preferences to avoid bias, the offline setting enforces a crucial tradeoff between noise and bias, making the choice of $\hat{\gamma}$ particularly critical.
Remark 4.6 (Discussions on $\alpha_{r}$ and $\alpha_{w}$).
The parameters $\alpha_{r}$ and $\alpha_{w}$ in Lemma˜4.2 are not fixed constants but are bounded within setting-dependent ranges. This design reflects the practical difficulty of determining each user’s exact neighbors during our theoretical analysis due to the uncertainty and noise in offline datasets, while still ensuring that the cardinalities of $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$ can be estimated reliably. These parameters therefore play a critical role in ensuring the generality of our results. Intuitively, by choosing $\alpha_{r}$ at the lower end of its range to conservatively approximate a subset of homogeneous neighbors and $\alpha_{w}$ at the upper end of its range to conservatively approximate a superset of heterogeneous neighbors, we can establish a concrete bound that remains parameter-fixed under uncertainty. We defer the refinement of these parameters, along with the derivation of a fixed parameter form of Theorem˜4.4, to Appendix B.1 due to space constraints.